OpenStudy (anonymous):

Hey I can't seem to find the formula for finding the area of a regular polygon with dies of 2cm

OpenStudy (anonymous):

Area = .5nsr where n is the number of sides, s is the length of each side, and r is the apothem, or the radius of the inscribed circle. Not 100%, but I believe r is \[r = \frac {1}{2} s*cot \frac {180}{n} \]

OpenStudy (anonymous):

The apothem is also \[r = \frac {1}{2} s * tan \frac {\pi (n-2)}{2n} \]

OpenStudy (anonymous):

Thanks